Consider that an object weighing 50 lb is dropped from a height of 1000ft with zero initial velocity. Assume that the air resistance is proportional to the velocity of the body. If the limiting velocity is known to be 200ft/sec, find the time it would take for an object to reach the ground.


1
Expert's answer
2021-11-30T16:01:45-0500

From the second Newton's Law

"mv'=mg-kv"

Given "m=(50\times 0.453592)kg=22.6796kg" and "g=32.1740ft\/s^2"


"v'+{k\over 22.6796}v=32.1740"


"v'=-{k\over 22.6796}(v-{729.6934504\over k})"


"{dv\over v-{729.6934504\over k}}=-{k\over 22.6796}dt"

We then integrate as follows


"v(t)={729.6934504\over k}-{729.6934504\over k}e^{-{kt\over 22.6796}}"


"v(t)\le200=\gt {729.6934504\over k}=200 \implies k=3.648467252"



"v(t)=200-200e^{-0.16087t}"


"v(t)=-h'(t)"


"h(t)=-\int vdt=-\int (200-200e^{-0.16087t})dt"


"=-200t-{200\over 0.16087}e^{-0.16087t}+c_2"


"h(0)=1000=-{200\over 0.16087}+c_2\implies c_2=2243.24"


"h(t)=-200t-1243.24e^{-0.16087t}+2243.24"


"h(t_1)=0=-200t_1-1243.24e^{-0.16087t_1}+2243.24"


And by solving this graphically, It will take 9.985 sec for the object to reach the ground